when the mother turned 46, her son 24. Presently, the mother is 5 times as old as her son. how old is the son now?
Answer by ankor@dixie-net.com(22740)

when the mother turned 46, her son 24.
m = s + 22
Presently, the mother is 5 times as old as her son.
m = 5s
how old is the son now?
replace m with s + 22
s + 22 = 5s
22 = 5s - s
22 = 4s
s = 22/4
s = 5.5 yr is son's age
then
5(5.5) = 27.5 is Mom's age